To attract more visitors, zoo authority announces 20% discount on every ticket which cost 29 paise. For this reason sale increases by 28% . Find the percentage of increase in the number of visitors.


1
Expert's answer
2021-02-24T07:49:29-0500

Let the number of visitors in beginning = 100

Total revenue gains = 0.29 * 100 = 29 Rs

after giving the discount new price of ticket "=29*\frac{80}{100}=(23.2)paise =0.232" Rs

New revenue "=29*\frac{128}{100}=37.12" Rs

If the number of visitors will be x then : "x*0.232=37.12"


"x=\frac{37.12}{0.232}=160"


percentage change in number of visitors "=\frac{160-100}{100}=60"


hence the % of increase in number of visitors = 60